As a manager of the XYZ Magical Supply Company, you were given the following demand schedule, what is the slope of the trend line ( if the equation is Qty = b + (m * Price); I am looking for the value of m). Pick the closest value.

Price

Qty

0

800

2

690

4

550

6

350

8

300

10

200
Answer #1

4= ar+b. a= (nEny) - (Ency) nexe) - (Ex)² n = 6. b = ({4{x²) - (Ex Exy) Qty. (nEx²)_(8X²) 800 70VA 650 OUT 600.767 550 500700